Suppose that price of labor is Rs.300 and the price of capital is Rs.400.what is the equation of the isocost line corresponding to a total cost of Rs.1000?
1
Expert's answer
2017-12-10T07:13:06-0500
If the price of labor is Rs.300 and the price of capital is Rs.400, then the isocost line corresponding to a total cost of Rs.1000 will connect points (0 units of labor; 2.5 units of capital) and (3.33; 0). So, the equation is K = 2.5 - 2.5/3.33*L, K = 2.5 - 0.75L.